# Phys2Real: Fusing VLM Priors with Interactive Online Adaptation for Uncertainty-Aware Sim-to-Real Manipulation

### v1

@@ Abstract
Learning robotic manipulation policies directly in the real world can be expensive and time-consuming. While reinforcement learning (RL) policies trained in simulation present a scalable alternative, effective sim-to-real transfer remains challenging, particularly for tasks that require precise dynamics. To address this, we propose Phys2Real, a real-to-sim-to-real RL pipeline that combines vision-language model (VLM)-inferred physical parameter estimates with interactive adaptation through uncertainty-aware fusion. Our approach consists of three core components: (1) high-fidelity geometric reconstruction with 3D Gaussian splatting, (2) VLM-inferred prior distributions over physical parameters, and (3) online physical parameter estimation from interaction data. Phys2Real conditions policies on interpretable physical parameters, refining VLM predictions with online estimates via ensemble-based uncertainty quantification. On planar pushing tasks of a T-block with varying center of mass (CoM) and a hammer with an off-center mass distribution, Phys2Real achieves substantial improvements over a domain randomization baseline: 100% vs 79% success rate for the bottom-weighted T-block, 57% vs 23% in the challenging top-weighted T-block, and 15% faster average task completion for hammer pushing. Ablation studies indicate that the combination of VLM and interaction information is essential for success. Project website: https://phys2real.github.io/
@@ I Introduction
Humans demonstrate remarkable exploration behaviors when manipulating novel objects in unseen settings. Initial judgments are formed about an object’s physical properties from visual appearance, and then these estimates are refined through interaction. This integration of perception and physical reasoning enables humans to adapt their manipulation strategies to specific instances of objects with potentially varying properties without extensive experience with each new object. Inspired by this observation, our approach seeks to provide robots with a similar ability to estimate and adapt to physical properties.
@@
We evaluate our approach on two planar pushing tasks: (1) T-block pushing with varying center of mass (CoM) achieved by placing a small metal weight at different locations of a 3D printed T-block, and (2) Hammer pushing with an off-center mass distribution. Our results demonstrate consistent improvements: Phys2Real achieves a success rate of 100% vs 79% for DR when the weight is at the bottom of the T-block, and 57.14% vs 23% for DR in the more challenging configuration with weight at the top of the T-block. Furthermore, on the hammer pushing task, it achieves 15% faster average task completion.
This work demonstrates that VLMs can provide interpretable, uncertainty-calibrated physical estimates for manipulation that can be combined with interaction history. Phys2Real is a step towards more general and adaptive robotic systems that learn from both perception and physical interaction.
@@ II-A Domain Randomization and System Identification
The sim-to-real gap remains a fundamental challenge for deploying policies learned in simulation to real-world environments. One class of approaches performs domain randomization (DR) [1, 4] to address this, randomizing simulation dynamics during training to develop policies robust to a range of environmental variations. Although DR-trained policies have shown success in several domains [5, 6, 7, 8], they often default to averaged behaviors that sacrifice performance for robustness, failing to adapt to object-specific variations. Our method trains simulation parameter-conditioned policies that are adapted online to specific conditions, rather than an unconditioned policy trained to be robust across all conditions.
@@ III Real-to-Sim-to-Real Policy Learning with Uncertainty-Aware Adaptation
Our method, Phys2Real, addresses the challenge of sim-to-real transfer by creating physically-informed digital twins and training adaptive manipulation policies. As illustrated in Fig. LABEL:pipeline, our approach consists of three stages: (1) real-to-sim reconstruction to create geometrically accurate simulation assets, (2) physics-parameter conditioned policy learning in simulation with uncertainty-aware adaptation, and (3) sim-to-real transfer that fuses VLM priors with estimates inferred from interaction data. Next, we describe each stage in detail.
Figure 1: Real-to-sim mesh reconstruction pipeline. Starting from a video of the object, we extract image frames and segment the target object using SAM-2. We then train a GSplat and extract a surface-aligned object-centric mesh using SuGaR [32]. Finally, we generate a clean, watertight mesh, resulting in a simulation-ready asset.
@@ III-A Real-to-Sim Scene Reconstruction
Figure 2: Phys2Real policy training. The policy and adaptation models are trained in three stages, inspired by RMA [2]. Phase 1: the policy is conditioned on ground truth physical properties (e.g., CoM) from simulation, Phase 1.5: fine-tune with noisy physical properties to build robustness to downstream noisy estimates from the fused VLM and adaptation estimate. Phase 2: train an ensemble of NN encoders that take in a history of observations and actions. The variance of the ensemble estimates provides an epistemic uncertainty (Eq. 1). Each encoder outputs a physical property estimate and an associated uncertainty, representing the model’s aleatoric uncertainty (Eq. 2). At test time, we fuse the adaptation estimates and VLM estimates via inverse-variance weighting (Eq. 4).
@@ III-B Physics-Conditioned Policy Learning
We train deep RL policies in simulation to manipulate the simulated objects. Motivated by evidence that policies with access to physical parameter information can improve performance over domain randomization approaches in locomotion [2], we explicitly condition the policy networks on physical parameter values.
For the parameter-conditioned policy to be deployable in the real world, we perform a second phase of training, during which the RL policy is frozen, and another neural network adaptation model learns to predict physical properties from historical state-action sequences. This approach is inspired by rapid motor adaptation (RMA) [2], which follows a similar two-phase learning paradigm: in the first phase, a physics encoder uses privileged information available only in simulation to produce a latent $z$ (representing physical properties), which is then input into the RL policy. In the second phase, history information is fed into an adaptation model, which regresses on the latent $z$ trained with the physics encoder from the first phase.
@@
    Phase 1.5 (optional): Since the policy at training time observes only ground truth physical parameter values, slightly incorrect VLM or physics encoder estimates at deployment time could lead to out-of-training-distribution states. We thus fine-tune the policy with noisy parameter estimates by applying random noise to each physical parameter during each episode. For our main experiments (Tables I-III), our policies were fine-tuned with a Gaussian noise of $\sigma=1.5\,\mathrm{cm}$.
@@
    Phase 2: We freeze the learned policy weights and train an ensemble of N=10{N=10} adaptation models that learn to predict physical parameters from a history of observations and actions over a sliding window. These models allow the policy to perform online adaptation during deployment. We obtain uncertainty estimates for the physical parameters using ensemble-based uncertainty quantification that captures both epistemic (model uncertainty) and aleatoric (data uncertainty) components.
@@ III-C Sim-to-Real Transfer with Physical Parameter Estimation
Figure 3: VLM priors for task-relevant physical parameters. We query a VLM (GPT-5 [36]) to provide an estimated CoM and uncertainty given an image of the object. For each of NN images, we repeat the query $M$ times. We then calculate the average VLM estimate $\bar{\theta}_{\textrm{vlm}}$ along with the average uncertainty $\bar{\sigma}_{\textrm{vlm}}$. This is fused with the RMA estimate using Eq. (4).
@@ III-C1 Obtaining $\theta_{\mathrm{vlm}}$ and $\sigma_{\mathrm{vlm}}$
We query GPT-5 [36] to estimate task-relevant physical parameters from images. For both the T-block and the hammer, we estimate the CoM along the vertical stem, which affects its rotational dynamics. As shown in Fig. 3, we capture $N$ images from different viewpoints and query the VLM $M$ times for each image. We use the prompt shown in Fig. 3 to obtain physical parameter and uncertainty estimates. Example VLM responses and estimation biases are provided in Appendix B.
The VLM estimate is computed as the aggregate mean across the prompts: $\theta_{\textrm{vlm}}=\frac{1}{N\times M}{\sum_{i=1}^{N}\sum_{j=1}^{M}\theta_{i,j}}$.
@@
TABLE I: T-block pushing with the weight at the top (placed $9.5\,\mathrm{cm}$ above the geometric center along the vertical axis; 21 trials per method with initial orientations varied every $45^{\circ}$). Arrows indicate whether higher ($\uparrow$) or lower ($\downarrow$) values are better. The privileged baseline is shaded in orange, and our method (Phys2Real) is highlighted in green with a star (*). The top two values in each column are bolded.
@@
Phys2Real* $\mathrm{CoM}=+4.0\,\mathrm{cm}$, $\boldsymbol{\sigma}=\pm 1.4\,\mathrm{cm}$] 57.14 2.60 $\pm$ 0.90 2.62 $\pm$ 1.73 39.58
@@
Phys2Real* $\mathrm{CoM}=+0.76\,\mathrm{cm}$, $\boldsymbol{\sigma}=\pm 1.55\,\mathrm{cm}$] 100.00 1.76 $\pm$ 0.54 4.73 $\pm$ 2.68 44.28
@@ III-C2 Obtaining $\theta_{\mathrm{rma}}$ and $\sigma_{\mathrm{rma}}$
$\sigma^{2}_{\textrm{epistemic}}=\frac{1}{N}\sum(\theta_{i}-\theta_{\mathrm{rma}})^{2}.$ (1)
@@
$\sigma^{2}_{\mathrm{aleatoric}}=\frac{1}{N}\sum\sigma^{2}_{i}.$ (2)
@@ III-C3 Fusing Estimates With Inverse-Variance Weighting
We combine physical property estimates from a VLM with estimates from RMA by weighting each source according to its uncertainty. This fusion mechanism allows the system to assign greater weight to VLM physical property estimates when the RMA interaction history is uncertain, and conversely, rely more on RMA when the VLM estimates are less certain. The physically interpretable nature of $\theta$ enables direct combination of VLM priors with online adaptation, unlike approaches that use learned latent representations.
@@
$\hat{\theta}=\frac{\theta_{\mathrm{vlm}}/\sigma_{\mathrm{vlm}}^{2}+\theta_{\mathrm{rma}}/\sigma_{\mathrm{rma}}^{2}}{1/\sigma_{\mathrm{vlm}}^{2}+1/\sigma_{\mathrm{rma}}^{2}}.$ (4)
where $\theta_{\mathrm{vlm}}$ and $\theta_{\mathrm{rma}}$ are estimates from the VLM and the ensemble of adaptation models, and $\sigma_{\mathrm{vlm}}$ and $\sigma_{\mathrm{rma}}$ are their respective uncertainties. $\hat{\theta}$ is used to condition the policy at test time.
@@ IV Experiments
Q4: Can Phys2Real be applied to real-world objects without available object models, and does policy performance still improve for them?
@@
  • T-block pushing: We vary the CoM along the vertical axis of the T-block by adding a small metal weight of 143 grams. We evaluate all policies using two different configurations based on where the metal weight is placed, which alters the CoM of the object along the vertical axis: (1) weight on the top, where the weight is placed at $9.5\,\mathrm{cm}$ above the geometric center of the T-block, and (2) weight on the bottom, where the weight is placed $6.5\,\mathrm{cm}$ below the geometric center. This makes the ground truth CoM $6.1\,\mathrm{cm}$ and $0.7\,\mathrm{cm}$ for the weights at top and bottom configurations, respectively. This task tests adaptation to varying object CoMs, which impacts the block’s rotational dynamics during pushing. We use a known T-block mesh (as in diffusion policy [37]) to isolate the effects of physical parameter estimation. We answer (Q1)-(Q3) with this task.
@@ IV-B1 Effect of Physical Parameter Estimation
The first two methods are baselines (DR and diffusion policy) that are physics parameter unaware. For DR, the location of the weight on the T-block is randomized in simulation from the bottom to the top of the block. We use diffusion policy [37] as a strong imitation learning baseline that is also unaware of object physics parameters. We train a state-based diffusion policy using motion capture data, matching our state-based RL policies. We collect $100$ demonstrations, varying the weight’s location on the vertical axis of the T-block and the T-block’s initial orientation in each episode, which is comparable to the domain randomization in simulation. We follow the implementation of  [37].
@@
The third method is a physics-conditioned policy with privileged information on the ground truth physics parameter.
@@
When we turn to the position error CDF in (Fig. 4b) for a more fine-grained understanding of position errors, we see that the Phys2Real curve rises sharply and up, indicating consistently low position error across nearly all trials. While diffusion policy has low error for around 40 percent of trials, the long tail is high. DR similarly has a long tail of high errors, showing that diffusion policy and DR easily become out of distribution, leading to high position errors.
When the weight is at the top of the block, shown in Table I, all policies struggle to achieve a high success rate. We hypothesize that the higher CoM causes more unpredictable dynamics. Although the privileged CoM-conditioned policy performs the best at 90.48%, showing that a more accurate physical estimation can indeed lead to better policy performance, Phys2Real is the second-best performer at 57.14%.
(a) Object and $E$ Trajectories. The end-effector (blue) and T-block (orange) paths are shown as the block moves $40\,\mathrm{cm}$ to the right from an initial orientation of $180^{\circ}$ (light blue outline) to a final orientation of $0^{\circ}$ (dark blue filled shape) at the target location (gold star). The robot achieves a final position error of $1.7\,\mathrm{cm}$.
(b) Parameter estimates over time. The VLM prior (red) provides an initial estimate ($\theta\approx 4.0\,\mathrm{cm}\pm 1.4\,\mathrm{cm}$).

### v2

@@ Abstract
Learning robotic manipulation policies directly in the real world can be expensive and time-consuming. While reinforcement learning (RL) policies trained in simulation present a scalable alternative, effective sim-to-real transfer remains challenging, particularly for tasks that require precise dynamics. To address this, we propose Phys2Real, a real-to-sim-to-real RL pipeline that combines vision-language model (VLM)-inferred physical parameter estimates with interactive adaptation through uncertainty-aware fusion. Our approach consists of three core components: (1) high-fidelity geometric reconstruction with 3D Gaussian splatting, (2) VLM-inferred prior distributions over physical parameters, and (3) online physical parameter estimation from interaction data. Phys2Real conditions policies on interpretable physical parameters, refining VLM predictions with online estimates via ensemble-based uncertainty quantification. On planar pushing tasks of a T-block with varying center of mass and a hammer with an off-center mass distribution, Phys2Real achieves substantial improvements over a domain randomization baseline: 100% vs 79% success rate for the bottom-weighted T-block, 57% vs 23% in the challenging top-weighted T-block, and 15% faster average task completion for hammer pushing. Ablation studies indicate that the combination of VLM and interaction information is essential for success. Project website: https://phys2real.github.io/
@@ I Introduction
Humans demonstrate remarkable exploration behaviors when manipulating novel objects in unseen settings. Initial judgments are formed about an object’s physical properties from visual appearance, and then these estimates are refined through interaction [6]. This integration of perception and physical reasoning enables humans to adapt their manipulation strategies to new instances of objects with potentially varying properties without extensive interaction. Inspired by this observation, our approach seeks to provide robots with a similar ability to estimate and adapt to physical properties.
@@
We evaluate our approach on two planar pushing tasks: (1) T-block pushing with varying center of mass (CoM) achieved by placing a small metal weight at different locations of a 3D printed T-block, and (2) Hammer pushing with an off-center mass distribution. Our results demonstrate consistent improvements: Phys2Real achieves a success rate of 100% vs. 79% for DR when the weight is at the bottom of the T-block, and 57.14% vs. 23% for DR in the more challenging configuration with weight at the top of the T-block. Furthermore, on the hammer pushing task, it achieves 15% faster average task completion.
This work demonstrates that VLMs can provide interpretable, uncertainty-calibrated physical estimates for manipulation that can be combined with interaction history. Phys2Real is a step towards more general, adaptive robotic systems that learn from perception and physical interaction.
@@ II-A Domain Randomization and System Identification
The sim-to-real gap remains a fundamental challenge for deploying policies learned in simulation to real-world environments. One class of approaches performs domain randomization (DR) [32, 25] to address this, randomizing simulation dynamics during training to develop policies robust to a range of environmental variations. Although DR-trained policies have shown success in several domains [30, 23, 7, 3], they often default to averaged behaviors that sacrifice performance for robustness, failing to adapt to object-specific variations. Our method trains simulation parameter-conditioned policies that adapt online to specific conditions, rather than an unconditioned policy trained to be robust across all conditions.
@@ III Real-to-Sim-to-Real Policy Learning with Uncertainty-Aware Adaptation
Our method, Phys2Real, addresses the challenge of sim-to-real transfer by creating physically-informed digital twins and training adaptive manipulation policies. As illustrated in Fig. LABEL:pipeline, our approach consists of three stages: (1) real-to-sim reconstruction to create geometrically accurate simulation assets, (2) physics-parameter conditioned policy learning in simulation with uncertainty-aware adaptation, and (3) sim-to-real transfer that fuses VLM priors with estimates inferred from interaction data.
Figure 1: Real-to-sim mesh reconstruction pipeline. From a video of the object, we extract image frames and segment the target object using SAM-2. We then train a GSplat and extract a surface-aligned object-centric mesh using SuGaR [9]. Finally, we generate a clean, watertight mesh, resulting in a simulation-ready asset.
@@ III-A Real-to-Sim Scene Reconstruction
Figure 2: Phys2Real policy training. The policy and adaptation models are trained in three stages, inspired by RMA [12]. Phase 1: the policy is conditioned on ground truth physical properties (e.g., CoM) from simulation, Phase 1.5: fine-tune with noisy physical properties to build robustness to downstream noisy estimates from the fused VLM and adaptation estimate. Phase 2: train an ensemble of $M$ encoders that take in a history of observations and actions. The variance of the ensemble estimates provides an epistemic uncertainty (Eq. 1). Each encoder outputs a physical property estimate and an associated uncertainty, representing the model’s aleatoric uncertainty (Eq. 2). At test time, we fuse the adaptation estimates and VLM estimates via inverse-variance weighting (Eq. 4).
@@ III-B Physics-Conditioned Policy Learning
We train deep RL policies in simulation to manipulate objects. Motivated by evidence that policies with access to physical parameter information can improve performance over domain randomization approaches in locomotion [12], we explicitly condition the policy networks on physical parameter values.
To enable real-world deployments, we perform a second phase of training, during which the RL policy is frozen, and another neural network adaptation model learns to predict physical properties from historical state-action sequences. This approach is inspired by rapid motor adaptation (RMA) [12], which follows a similar two-phase learning paradigm: in the first phase, a physics encoder uses privileged information available only in simulation to produce a latent $z$ (representing physical properties), which is then input into the RL policy. In the second phase, history information is fed into an adaptation model, which regresses on the latent $z$ trained with the physics encoder from the first phase.
@@
    Phase 1.5 (optional): Since the policy at training time observes only ground truth physical parameter values, slightly incorrect VLM or physics encoder estimates at deployment time could lead to out-of-training-distribution states. We thus fine-tune the policy with noisy parameter estimates by applying random noise to each physical parameter during each episode. Our policies were fine-tuned with a Gaussian noise of $\sigma=1.5\,\mathrm{cm}$.
@@
    Phase 2: We freeze the learned policy weights and train an ensemble of M=10{M=10} adaptation models that learn to predict physical parameters from a history of observations and actions over a sliding window of H=10{H=10}. These models allow the policy to perform online adaptation during deployment. We obtain uncertainty estimates for the physical parameters using ensemble-based uncertainty quantification that captures both epistemic (model uncertainty) and aleatoric (data uncertainty) components.
@@ III-C Sim-to-Real Transfer with Physical Parameter Estimation
Figure 3: VLM priors for task-relevant physical parameters. We query a VLM (GPT-5 [24]) to provide an estimated CoM and uncertainty given an image of the object. For each of $V$ images, we repeat the query $Q$ times. We then calculate the average VLM estimate $\bar{\theta}_{\textrm{vlm}}$ along with the average uncertainty $\bar{\sigma}_{\textrm{vlm}}$. This is fused with the RMA estimate using Eq. (4).
@@ III-C1 Obtaining $\theta_{\mathrm{vlm}}$ and $\sigma_{\mathrm{vlm}}$
We query GPT-5 [24] to estimate task-relevant physical parameters from images. For both the T-block and the hammer, we estimate the CoM along the vertical stem, which affects its rotational dynamics. As shown in Fig. 3, we capture $V$ images from different viewpoints and query the VLM $Q$ times for each image. We use the prompt shown in Fig. 3 to obtain physical parameter and uncertainty estimates.
Example VLM responses and estimation biases are provided in Appendix B.
The VLM estimate is computed as the aggregate mean across the prompts: $\theta_{\textrm{vlm}}=\frac{1}{V\times Q}{\sum_{i=1}^{V}\sum_{j=1}^{Q}\theta_{i,j}}$.
@@
TABLE I: T-block pushing with the weight at the top (placed $9.5\,\mathrm{cm}$ above the geometric center along the vertical axis; 21 trials per method with initial orientations varied every $45^{\circ}$). Arrows indicate whether higher ($\uparrow$) or lower ($\downarrow$) values are better, with the top-two values shown in bold. The privileged baseline is shaded in orange, and our method (Phys2Real) is highlighted in green with a star (*).
@@
Phys2Real* $\mathrm{CoM}=+4.0\,\mathrm{cm}$, $\boldsymbol{\sigma}=1.4\,\mathrm{cm}$] 57.14 2.60 $\pm$ 0.90 2.62 $\pm$ 1.73 39.58
@@
Phys2Real* $\mathrm{CoM}=+0.76\,\mathrm{cm}$, $\boldsymbol{\sigma}=1.55\,\mathrm{cm}$] 100.00 1.76 $\pm$ 0.54 4.73 $\pm$ 2.68 44.28
@@ III-C2 Obtaining $\theta_{\mathrm{rma}}$ and $\sigma_{\mathrm{rma}}$
$\sigma^{2}_{\textrm{epistemic}}=\frac{1}{M}\sum(\theta_{i}-\theta_{\mathrm{rma}})^{2}.$ (1)
@@
$\sigma^{2}_{\mathrm{aleatoric}}=\frac{1}{M}\sum\sigma^{2}_{i}.$ (2)
@@ III-C3 Fusing Estimates With Inverse-Variance Weighting
We combine physical property estimates from a VLM with estimates from RMA using uncertainty-based weights.
This fusion mechanism allows the system to assign greater weight to VLM physical property estimates when the RMA interaction history is uncertain, and conversely, rely more on RMA when the VLM estimates are less certain. The physically interpretable nature of $\theta$ enables direct combination of VLM priors with online adaptation, unlike approaches that use learned latent representations.
@@
$\hat{\theta}=\frac{\theta_{\mathrm{vlm}}/\sigma_{\mathrm{vlm}}^{2}+\theta_{\mathrm{rma}}/\sigma_{\mathrm{rma}}^{2}}{1/\sigma_{\mathrm{vlm}}^{2}+1/\sigma_{\mathrm{rma}}^{2}},$ (4)
where $\theta_{\mathrm{vlm}}$ and $\theta_{\mathrm{rma}}$ are estimates from the VLM and the ensemble of adaptation models, and $\sigma_{\mathrm{vlm}}$ and $\sigma_{\mathrm{rma}}$ are their respective uncertainties. $\hat{\theta}$ is used to condition the policy at test time. Under the assumption that the estimates are unbiased with independent noise and known variances, this estimator is theoretically justified as the Best Linear Unbiased Estimator (BLUE), which minimizes the variance of the fused estimate.
@@ IV Experiments
Q4: Can Phys2Real improve policy performance when applied to real-world objects without available object models?
@@
  • T-block pushing: To answer (Q1)-(Q3), we vary the CoM along the vertical axis of the T-block by adding a small metal weight of 143 grams. We evaluate all policies using two different configurations based on where the metal weight is placed, which alters the CoM of the object along the vertical axis: (1) weight on the top, where the weight is placed at $9.5\,\mathrm{cm}$ above the geometric center of the T-block, and (2) weight on the bottom, where the weight is placed $6.5\,\mathrm{cm}$ below the geometric center. This makes the ground truth CoM $6.1\,\mathrm{cm}$ and $0.7\,\mathrm{cm}$ for the weights at top and bottom configurations, respectively. This task tests adaptation to varying object CoMs, which impacts the block’s rotational dynamics during pushing. We use a known T-block mesh (as in DP [4]) to isolate the effects of physical parameter estimation.
@@ IV-B1 Effect of Physical Parameter Estimation
The first two methods are baselines, DR and diffusion policy (DP), that are physics parameter unaware. For DR, the location of the weight on the T-block is randomized in simulation from the bottom to the top of the block. We use DP [4] as a strong imitation learning baseline that is also unaware of object physics parameters. We train a state-based DP using motion capture data, matching our state-based RL policies. We collect $100$ demonstrations, varying the weight’s location on the vertical axis of the T-block and the T-block’s initial orientation in each episode, which is comparable to the domain randomization in simulation. We follow the implementation of  [4].
@@
The third method is a physics-conditioned policy with privileged information on the ground truth physics parameter, computed from the known T-block geometry assuming uniform density and the metal weight’s measured mass. This baseline represents the oracle upper bound on performance.
@@
When we turn to the position error CDF in (Fig. 4b) for a more fine-grained understanding of position errors, we see that the Phys2Real curve rises sharply and up, indicating consistently low position error across nearly all trials. While DP has low error for around 40 percent of trials, the long tail is high. DR similarly has a long tail of high errors, showing that DP and DR easily become out of distribution, leading to high position errors. When the weight is at the top of the block, shown in Table I, all policies struggle to achieve a high success rate, except the privileged CoM-conditioned policy with a 90.48% success rate. We hypothesize that the higher CoM causes more unpredictable dynamics. Even in this case, Phys2Real is the second-best performer at 57.14%.
(a) Object and end-effector trajectories. The end-effector (blue) and T-block (orange) paths are shown as the block moves $40\,\mathrm{cm}$ to the right from an initial orientation of $180^{\circ}$ (light blue outline) to a final orientation of $0^{\circ}$ (dark blue filled shape) at the target location (gold star). The robot achieves a final position error of $1.7\,\mathrm{cm}$.
(b) Parameter estimates over time. The VLM prior (red) provides an initial estimate ($\theta_{\mathrm{vlm}}=4.0\,\mathrm{cm}$, $\sigma_{\mathrm{vlm}}=1.4\,\mathrm{cm}$).